OpenStudy (anonymous):

Complete the statement: if |2x-1|<2 and |3-x|<5 then: Can you please explain each step please? I have a a lot more of these types of questions and I dunno how to do them.

OpenStudy (anonymous):

if \(|2x-1|<2\) then \[-2<2x-1<2\\ -1<2x<3\\ -\frac{1}{2}<x<\frac{3}{2}\]

OpenStudy (anonymous):

I forgot to say but there are answer choices: A. \[\left| x-\frac{ 1 }{ 2 } \right|<1\] B. \[\left| 3x-4 \right|<7\] C. \[\left| x-\frac{ 1 }{ 2 } \right|>1\] D.\[\left| x-4 \right|>1\] E.\[\left| x-4 \right|<1\]

OpenStudy (aum):

\(|2x-1| \lt 2\) divide by 2: \(\Large \frac{|2x-1|}{2} \lt \frac 22\) \(\Large |x-\frac 12|\lt 1\)

OpenStudy (anonymous):

But how about the second inequality? What do I do with |3-x|<5 ?